OpenStudy (galacticplunder):

9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (galacticplunder):

@SolomonZelman @jigglypuff314 @imammalik_806 @BPDlkeme234

OpenStudy (galacticplunder):

Guys!

OpenStudy (danjs):

that is 2 variables in one equation...there will be infinitely many solution points(x,y) on that line

OpenStudy (danjs):

what do you do now, set notation or

OpenStudy (danjs):

solutions - {all (x,y) such that, 9x - y = -2}

OpenStudy (danjs):

or, maybe, solve for y, y = 9x + 2 solution is all points (x , y) = (x , 9x + 2)

OpenStudy (galacticplunder):

?

OpenStudy (galacticplunder):

Sorry I have no idea what you're talking about

OpenStudy (danjs):

Do you see, the equation is a line?

OpenStudy (galacticplunder):

Yeah

OpenStudy (danjs):

9x - y =- 2 or y = 9x + 2

OpenStudy (galacticplunder):

Umm... ok. How did you do that?

OpenStudy (danjs):

i just rearranged it to the y = mx + b form

OpenStudy (danjs):

you don't really need to

OpenStudy (galacticplunder):

Um, ok

OpenStudy (galacticplunder):

So i mean how do i explain it?

OpenStudy (galacticplunder):

What I do for part A

OpenStudy (danjs):

i guess you can draw the graph, slope 9 and y intercept 2, it is slanted then explain the solutions are points (x,y) on that line, infinitely many of them. Maybe say since y=9x+2, each point will be in the form (x,y) = (x , 9x + 2)

OpenStudy (galacticplunder):

um

OpenStudy (galacticplunder):

So I say that you graph it?

OpenStudy (danjs):

The solutions are all points (x,y) such that y = 9x+2

OpenStudy (galacticplunder):

Oh ok

OpenStudy (danjs):

(x , y) = (x, 9x+2)

OpenStudy (galacticplunder):

ok

OpenStudy (galacticplunder):

What about part B?

OpenStudy (danjs):

pick any 3 values for x, and show the resulting points (x,y)

OpenStudy (galacticplunder):

ok

OpenStudy (galacticplunder):

ok i get it

OpenStudy (danjs):

IF x=0 --- point (0,2)

OpenStudy (galacticplunder):

cool

OpenStudy (galacticplunder):

Ok so i have that

OpenStudy (galacticplunder):

What about part c?

OpenStudy (danjs):

inconsistent system means there will be no solutions... to show that here...

OpenStudy (danjs):

you want any parallel line, that is not the same line

OpenStudy (danjs):

2 parallel lines never intersect, and thus have no solutions

OpenStudy (galacticplunder):

ok so I would change the y intercept?

OpenStudy (danjs):

in the form given a*x + b*y = c another general d*x + e*y = f a simple way to make a parallel line from that form, is to take a number multiple of the coefficients on x and y - ---

OpenStudy (danjs):

or yeah just change the y intercept in that form

OpenStudy (galacticplunder):

Thanks! Could you help me with another

OpenStudy (galacticplunder):

A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (danjs):

ok

OpenStudy (danjs):

When you are solving systems of equations - it is legal to -multiply any equation by a constant number, it remains the same line. -add equations together -change order of equations

OpenStudy (danjs):

So for that example, you just follow what it tells you to do... 8x + 5y = 9 ---first equation 6x + 4y = 8 ---2 times second equation ----------- + 14x + 9y = 17 --- new first equation

OpenStudy (galacticplunder):

i dont understand

OpenStudy (galacticplunder):

-_-

OpenStudy (galacticplunder):

oh wait

OpenStudy (galacticplunder):

nevermind got it thank you

OpenStudy (galacticplunder):

what about part b

OpenStudy (galacticplunder):

SOMEONE HELP